1. Determine whether the following is a set or not a set.


a. The list of course offering of UPHR-Molino Campus. -SET

b. The elected barangay officials of Bacoor City. -SET

c. The collection of intelligent students of College Department. -NOT A SET


2. List the elements of the following sets.


a. A = {x/x is a letter in the word mathematics}

b. B = {x/x is a positive integer, 3 ≤ x ≤ 8}

c. C = {x/x = 2n + 3, n is a positive integer}


3. Given U = {x/x is the set of letters in the English alphabet;


A = {a, e, i, o, u ,

B = {x/x is the set of consonant letters} ; and

C = { a, b, c, d, e}.



Tell whether the statement is true or false.


a. B ⊂ U

b. C ⊂ B

c. A ⊆ B

d. B = A


4. Determine the cardinal number sets in given No. 3 such as


a. universal set (U)

b. set A

c. set B

d. set C


1
Expert's answer
2021-10-07T14:01:35-0400

1.

A set is a collection of elements.

a. set

b. set

c. set


2.

a. "A=\{m,a,t,h,e,i,c,s\}"

b. "B=\{3,4,5,6,7,9\}"

c. "C=\{5,7,9,11,...\}"

"2\cdot 1+3=5,\ 2\cdot 2+3=7,\ 2\cdot 3+3=9" etc.


3.

a. true

b. false (a,e are not consonant letters)

c. false (a, e, i, o, u are not consonant letters)

d. false (a, e, i, o, u are not consonant letters)


4.

a. "n(U)=26"

b. "n(A)=5"

c. "n(B)=21"

b. "n(C)=5"